OpenStudy (anonymous):

Solve the equation. Show your work. 45 = 3b + 69

OpenStudy (texaschic101):

are you solving for b ?

OpenStudy (anonymous):

yes

OpenStudy (nurali):

45 = 3b + 69 3b=45-69 3b=-24 b=-24/3 b=-8

OpenStudy (texaschic101):

subtract 69 from both sides then divide both sides by 3

OpenStudy (anonymous):

thanks!

OpenStudy (nurali):

My Pleasure.

OpenStudy (anonymous):

\[45=3b +69\]\[45-69 = 3b\]\[-24=3b\]\[\frac{ -24 }{ 3}= b\]\[-8 = b\]

OpenStudy (anonymous):

Dammit lol, I make it look nice lol

OpenStudy (nurali):

Welcome to Openstudy @lailastoppler